Question Help * nd of Chapter 5.2 According to a study by the U.S. Centers for Disease Control and Prevention, the price elasticity of demand for cigarettes is 0.25. Americans purchase about 480 billion cigarettes each year Source: Response to Increases in Cigarette Prices by Race/Ethnicity, Income, and Age Groups-United States, 1976 -1993, Morbidity and Mortality Weekly Report, July 31, 1998 a. If the federal tax on cigarettes were increased enough to raise the price of cigarettes by 60 percent, the quantity of cigarettes demanded would by billion per year (enter your numeric response rounded to the nearest integer). decrease increase
Answer #1

Answer

Formula :

Price elasticity of demand = % change in quantity demand / % change in price----------------(1)

It is given that, Price elasticity of demand = -0.25.

Also it is given that because of federal tax , Price of cigarettes increased by 60%

Hence, % change in Price = 60

Using (1) and given information we get

Price elasticity of demand = % change in quantity demand / % change in price

=> -0.25 = % change in quantity demand / 60

=> % change in quantity demand = -0.25*60 = -15(this negative sign suggest that quantity demanded demand decreases).

Hence ,Quantity demanded decreases by 15%.

Initial quantity demand = 480 billion and Now quantity demand decreased by 15%.

Thus, quantity demand will decrease by 15% of 480 billion = (15/100)*480 billion = 72 billion.

Hence, (a) If the federal tax on cigarettes were increased enough to raise the price of cigarettes by 60 percent, the quantity of cigarettes demanded would decrease by 72 billion.